# D-SCo: Dual-Stream Conditional Diffusion for Monocular Hand-Held Object Reconstruction

### v3

@@ Abstract
Reconstructing hand-held objects from a single RGB image is a challenging task in computer vision. In contrast to prior works that utilize deterministic modeling paradigms, we employ a point cloud denoising diffusion model to account for the probabilistic nature of this problem. In the core, we introduce centroid-fixed Dual-Stream Conditional diffusion for monocular hand-held object reconstruction (D-SCo), tackling two predominant challenges. First, to avoid the object centroid from deviating, we utilize a novel hand-constrained centroid fixing paradigm, enhancing the stability of diffusion and reverse processes and the precision of feature projection. Second, we introduce a dual-stream denoiser to semantically and geometrically model hand-object interactions with a novel unified hand-object semantic embedding, enhancing the reconstruction performance of the hand-occluded region of the object. Experiments on the synthetic ObMan dataset and three real-world datasets HO3D, MOW and DexYCB demonstrate that our approach can surpass all other state-of-the-art methods. Codes will be released.
@@
Point Cloud Denoising Diffusion Model Hand-Held Object Reconstruction
@@ 1 Introduction
In essence, most existing works either rely on Signed Distance Fields (SDFs) [37, 78] or Directed Distance Fields (DDFs) [80] to represent object shapes. Despite being effective under favorable conditions, such techniques tend to result in over-smoothed and undetailed reconstruction [52, 11]. Moreover, prior work [31, 37, 78, 80] typically utilize a deterministic modeling paradigm, making it difficult to reason about the uncertainties introduced by hand- or self-occlusion. Recently, probabilistic point cloud denoising diffusion models [42, 83, 44] have shown to be effective for the task of single-view object reconstruction. Compared to works employing surface-based representations like SDFs and DDFs, diffusion-driven methods for reconstructing point clouds enjoy better capabilities in overcoming artifacts such as noise, sparsity, and irregularities [11]. Moreover, the probabilistic nature of denoising diffusion models is particularly beneficial for modeling uncertainties and underconstrained problems.
@@
To solve the aforementioned problems, we present D-SCo, a centroid-fixed dual-stream conditional point cloud denoising diffusion model for single-view hand-held object reconstruction. As shown in Fig. 1, we compare our D-SCo with naive diffusion models. First, a hand-constrained centroid fixing scheme is proposed to ensure that the centroid of the partially denoised point cloud does not diverge during the diffusion as well as the reverse process. In particular, we use a small and efficient neural network to estimate the hand-constrained object centroid and use it as a guide during the reverse process. Hence, the model only needs to consider the simpler shape diffusion task rather than having to simultaneously account for shape and position. Second, to best leverage the hand-object interaction prior, we introduce a dual-stream denoiser, which individually processes semantic and geometric priors. In particular, we utilize a unified hand-object semantic embedding to compensate for hand occlusion in the 2D feature extraction stage.
Our contributions can be summarized as follows.
@@
  • We further propose a dual-stream denoiser to semantically and geometrically model hand-object interactions. Our novel unified hand-object semantic embedding serves as a strong prior for reconstructing occluded regions of the object.
@@ 2 Related Work
[…]
@@ 3 Method
In the following sections, we first introduce the fundamentals of conditional point cloud denoising diffusion models (Sec. 3.1). We then explain our centroid-fixed conditional diffusion (Sec. 3.2) before diving into our proposed dual-stream conditional denoiser (Sec. 3.3).
@@ 3.2 Hand-Constrained Centroid-Fixed Conditional Diffusion
$\epsilon\sim\mathcal{N}(\mathbf{0},$ $~{}\mathbf{I}),\epsilon\leftarrow\epsilon-\bar{\epsilon}.$
Noteworthy, by keeping the object centroid unchanged, the misalignment error within the semantic feature projection due to centroid movements could also be alleviated. Thus, the training behavior becomes more stable as well.
@@ 3.3 Dual-Stream Conditional Point Cloud Denoising
Hence, we instead employ a dual-stream denoiser to separately process $X_{t}^{HO}$ and $X_{t}^{A}$. As shown in Fig. 2, given $z_{1}=X_{t}^{HO}$ and $z_{2}=X_{t}^{A}$, we first feed the object and hand points along with their corresponding semantic embedding $[X_{t},z_{1}]$ to one branch of the dual-stream denoiser $f_{\theta}^{1}$ and obtain $\mathcal{F}_{\theta}^{1}\in\mathbb{R}^{(N+N_{H})\times S}$ as feature representation guided by the semantic prior, with $S$ denoting the number of latent feature channels. Similarly, we feed the object point cloud along with its corresponding geometric embedding $[X_{t},z_{2}]$ to $f_{\theta}^{2}$ using the identical architecture to obtain $\mathcal{F}_{\theta}^{2}\in\mathbb{R}^{N\times S}$ as our geometric feature. The final noise is then predicted from the concatenation of the semantic and geometric features with $\epsilon_{\theta}=g_{\theta}([\mathcal{F}_{\theta}^{1},\mathcal{F}_{\theta}^{2}])\in\mathbb{R}^{N\times 3}$, where $g_{\theta}$ consists of stacked Multilayer Perceptrons (MLPs).
@@ 3.4 Training Objectives
where $\eta_{1}$ is a hyperparameter that controls the strength of the projective 2D regularization term.
@@ 4.1 Datasets and Setup
Evaluation Metrics. Following [78, 80], we report the Chamfer Distance (CD) in mm and F-score at thresholds of 5mm (F-5) and 10mm (F-10) to compare with the state-of-the-art.
Implementation Details are provided in supplementary material.
@@ 4.2 Evaluation on ObMan Dataset
We first present quantitative results on the large-scale synthetic ObMan dataset in Tab. 1 (a). As can be seen, our approach surpasses all state-of-the-art methods by a large margin for F-score and Chamfer Distance. Specifically, compared with the current best-performing method DDF-HO [80], we achieve a relative improvement of 10.9 % and 20.9 % for the F-5 and F-10 metrics, respectively. Furthermore, we can reduce the CD by 21.4 % compared with DDF-HO and 89.2 % compared with iHOI, demonstrating that our reconstructed shape possesses significantly fewer outliers.
@@ 4.3 Evaluation on Real-World Datasets
Zero-shot transfer to HO3D and MOW. To further evaluate our zero-shot transfer abilities, we also directly apply our model, trained on the ObMan dataset, to HO3D and MOW without conducting any additional finetuning. The results in Tab. 2 (bottom) show that our method can again achieve a remarkable improvement on HO3D (13 % and 14 % in F-5 and F-10) and MOW (71 % and 89 % in F-5 and F-10), demonstrating the strength of D-SCo in synthetic-to-real generalization.
@@ 4.4 Ablation Studies
Figure 4: Qualitative results on HO3D [28] (top) and MOW [4] (middle and bottom) datasets. For each method and ground truth, we show the reconstruction results in the camera view (column 1) and a novel view (column 2).
@@ F-5 $\uparrow$ F-10 $\uparrow$ CD $\downarrow$ F-5 $\uparrow$ F-10 $\uparrow$ CD $\downarrow$
The robustness against hand pose prediction quality is further discussed in supplementary material.
@@ 5 Conclusion
In this paper, we present D-SCo, a novel centroid-fixed dual-stream conditional diffusion model for single-view hand-held object reconstruction. D-SCo does not require any object templates, category priors, or depth information, and exhibits superior performance in modeling the uncertainties induced by hand- and self-occlusion. In the core, we propose a hand-constrained centroid-fixing paradigm, utilizing the estimated hand vertices to prevent the centroid of the partially denoised point cloud from diverging during diffusion and reverse processes. Further, a dual-stream denoiser is introduced to semantically and geometrically model hand-object interaction, with a novel unified hand-object semantic embedding enhancing the robustness against occlusion. Our experiments demonstrate that our approach is able to surpass existing methods in both synthetic and real-world scenarios.

### v4

@@ Abstract
Reconstructing hand-held objects from a single RGB image is a challenging task in computer vision. In contrast to prior works that utilize deterministic modeling paradigms, we employ a point cloud denoising diffusion model to account for the probabilistic nature of this problem. In the core, we introduce centroid-fixed Dual-Stream Conditional diffusion for monocular hand-held object reconstruction (D-SCo), tackling two predominant challenges. First, to avoid the object centroid from deviating, we utilize a novel hand-constrained centroid fixing paradigm, enhancing the stability of diffusion and reverse processes and the precision of feature projection. Second, we introduce a dual-stream denoiser to semantically and geometrically model hand-object interactions with a novel unified hand-object semantic embedding, enhancing the reconstruction performance of the hand-occluded region of the object. Experiments on the synthetic ObMan dataset and three real-world datasets HO3D, MOW and DexYCB demonstrate that our approach can surpass all other state-of-the-art methods.
@@
Point Cloud Diffusion Hand-Held Object Reconstruction
@@ 1 Introduction
In essence, most existing works either rely on Signed Distance Fields (SDFs) [37, 78] or Directed Distance Fields (DDFs) [80] to represent object shapes. Despite being effective under favorable conditions, such techniques generally tend to result in over-smoothed and undetailed reconstruction [52, 11]. Moreover, prior work [31, 37, 78, 80] typically utilize a deterministic modeling paradigm, making it difficult to reason about the uncertainties introduced by hand- or self-occlusion. Recently, probabilistic point cloud denoising diffusion models [42, 83, 44] have shown to be effective for the task of single-view object reconstruction. Compared to works employing surface-based representations like SDFs and DDFs, diffusion-driven methods for reconstructing point clouds enjoy better capabilities in overcoming artifacts such as noise, sparsity, and irregularities [11]. Moreover, the probabilistic nature of denoising diffusion models is particularly beneficial for modeling uncertainties and underconstrained problems.
@@
To solve the aforementioned problems, we present D-SCo, a centroid-fixed dual-stream conditional point cloud denoising diffusion model for single-view hand-held object reconstruction. As shown in Fig. 1, we compare our D-SCo with naive diffusion models. First, a hand-constrained centroid fixing scheme is proposed to ensure that the centroid of the partially denoised point cloud does not diverge during the diffusion as well as the reverse process. In particular, we use a small and efficient neural network to estimate the hand-constrained object centroid and use it as a guide during the reverse process. Hence, the model only needs to consider the simpler shape diffusion task rather than having to simultaneously account for shape and position. Second, to best leverage the hand-object interaction prior, we introduce a dual-stream denoiser, which individually processes semantic and geometric priors. In particular, we utilize a unified hand-object semantic embedding to compensate for hand occlusion in the 2D feature extraction stage. Our contributions can be summarized as follows.
@@
  • We propose a dual-stream denoiser to semantically and geometrically model hand-object interactions. Our novel unified hand-object semantic embedding serves as a strong prior for reconstructing occluded regions of the object.
@@ 2 Related Work
Diffusion Models for 3D Reconstruction. Denoising diffusion models (DDMs) [33, 49] have recently attracted increasing attention in 3D reconstruction. While Luo et al. [43] and LION [79] employ latent variables for the diffusion process in point cloud generation, PVD [83] directly applies DDMs to point clouds, leading to a unified framework for both unconditional 3D generation and conditional completion of partial shapes. Built upon PVD, PC2 [44] conducts single-view reconstruction by conditioning on projected image features. In this work, we employ point cloud diffusion models [83, 44] for hand-held object reconstruction, as we observed that these models are more robust towards producing fragmented or distorted surfaces under ambiguous views than SDF-driven approaches [78].
@@ 3.2 Hand-Constrained Centroid-Fixed Conditional Diffusion
$\epsilon\sim\mathcal{N}(\mathbf{0},$ $~{}\mathbf{I}),\epsilon\leftarrow\epsilon-\bar{\epsilon},$
where $\bar{X_{0}}$ denotes the centroid of $X_{0}$. Noteworthy, by keeping the object centroid unchanged, the misalignment error within the semantic feature projection due to centroid movements could also be alleviated. Thus, the training behavior becomes more stable as well.
@@ 3.3 Dual-Stream Conditional Point Cloud Denoising
Hence, we instead employ a dual-stream denoiser to separately process $X_{t}^{HO}$ and $X_{t}^{A}$. As shown in Fig. 2, given $z_{1}=X_{t}^{HO}$ and $z_{2}=X_{t}^{A}$, we first feed the object and hand points along with their corresponding semantic embedding $[X_{t},z_{1}]$ to one branch of the dual-stream denoiser $f_{\theta}^{1}$ and obtain $\mathcal{F}_{\theta}^{1}\in\mathbb{R}^{(N+N_{H})\times S}$ as feature representation guided by the semantic prior, with $S$ denoting the number of latent feature channels. Similarly, we feed the object point cloud along with its corresponding geometric embedding $[X_{t},z_{2}]$ to $f_{\theta}^{2}$ using the identical architecture to obtain $\mathcal{F}_{\theta}^{2}\in\mathbb{R}^{N\times S}$ as our geometric feature. The final noise is then predicted from the concatenation of the semantic and geometric features with $\epsilon_{\theta}=g_{\theta}([\mathcal{F}_{\theta}^{1},\mathcal{F}_{\theta}^{2}])\in\mathbb{R}^{N\times 3}$, where $g_{\theta}$ consists of stacked MLPs.
@@ 3.4 Training Objectives
where the weight $\eta_{1}$ controls the strength of the projective 2D regularization.
@@ 4.1 Datasets and Setup
Metrics. Following [78, 80], we report the Chamfer Distance (CD) in mm and F-score at thresholds of 5mm (F-5) and 10mm (F-10).
@@ 4.2 Evaluation on ObMan Dataset
We first present quantitative results on the large-scale synthetic ObMan dataset in Tab. 1 (a). As shown, our approach surpasses all state-of-the-art methods by a large margin for F-score and Chamfer Distance. Specifically, compared with the current best-performing method DDF-HO [80], we achieve a relative improvement of 10.9 % and 20.9 % for F-5 and F-10 metrics, respectively. Furthermore, we reduce the CD by 21.4 % compared with DDF-HO and 89.2 % compared with iHOI, demonstrating that our reconstructed shape possesses significantly fewer outliers.
@@ 4.3 Evaluation on Real-World Datasets
Zero-shot transfer to HO3D and MOW. To further evaluate our zero-shot transfer abilities, we also directly apply our model, trained on ObMan, to HO3D and MOW without conducting any additional finetuning. The results in Tab. 2 (bottom) show that our method again achieves a remarkable improvement on HO3D (13 % and 14 % in F-5 and F-10) and MOW (71 % and 89 % in F-5 and F-10), demonstrating D-SCo’s strength in synthetic-to-real generalization.
@@ 4.4 Ablation Studies
Figure 4: Qualitative results on HO3D [28] (top) and MOW [4] (middle and bottom) datasets. For each method and ground truth, we show the reconstruction results in the camera view (column 1) and a novel view (column 2). We also show the unoccluded objects for our method and ground truth.
@@ 5 Conclusion
This paper presents D-SCo, a novel centroid-fixed dual-stream conditional diffusion model for single-view hand-held object reconstruction. D-SCo does not require any object templates, category priors, or depth information, and excels at modeling uncertainties induced by hand- and self-occlusion. In the core, we propose a hand-constrained centroid-fixing paradigm, utilizing estimated hand vertices to prevent the centroid of the partially denoised point cloud from diverging during diffusion and reverse processes. Further, a dual-stream denoiser is introduced to semantically and geometrically model hand-object interaction, with a novel unified hand-object semantic embedding enhancing the robustness against occlusion.
Acknowledgements
This work was supported in part by the China Scholarship Council and in part by the Shuimu-Zhiyuan Tsinghua Scholar Program.